for a daily lottery, a person selects a three-digit number. If the person plays for $1, she can win $500. Find the expectation. In the same daily lottery, if a person boxes the number 123, she will win $80. Find the expectation if the number 123 is played for $1 and boxed
Answer by edjones(8007)

a)
The probability that any single number will be the lucky DISABLED_event_one= 1/10^3 = 1/1000
E=ap Where a is the amount and p the probability.
E = 500 * 1/1000 = $0.50
.
b)
123 box = 3! = 6 chances to win
E = 80 * 6/1000 = 480/1000 = $0.48
.
Ed